A spherical snowball is melting in such a way
that its diameter is decreasing at rate of 0.1 cm/min. At what
rate is the volume of the snowball decreasing when the diameter
is 13 cm. (Note the answer is a positive number)
1
Expert's answer
2020-03-10T12:13:27-0400

Given

Diameter if sphere is decreasing at the rate of 0.1 cm/min

"i.e.\:\frac{dD}{dt}=-0.1\:\frac{cm}{min}" (negative sign means diameter is decreasing)

We need to find rate of volume of snowball decreasing with respect to time "(\frac{dV}{dt})" when diameter (D) is 13 cm.

we know volume of Sphere is

"V=\frac{4}{3}\pi r^3"

we also know radius is equal to half of diameter

Substituting "r=\frac{D}{2}" in above equation

"V=\frac{4}{3}\pi \left(\frac{D}{2}\right)^3"

"V=\frac{\pi }{6} D^3"

Differentiating with respect to time

"\frac{dV}{dt}=\frac{\pi }{6}\cdot 3D^2\cdot \frac{dD}{dt}"

substituting values

"\frac{dV}{dt}=\frac{\pi }{6}\cdot 3\left(13\:cm\right)^2\cdot \left(-0.1\:\frac{cm}{min}\right)"

"\frac{dV}{dt}=-\frac{169\pi }{20}\:\frac{cm^3}{min}\approx -26.546\:\frac{cm^3}{min}"

Negative sign indicate volume is decreasing with respect to time.


Volume is decreasing at the rate of "\frac{169\pi }{20}" "cm^3\/min" or 26.546 "cm^3\/min"


Note:

1) Round answer as per requirements or write in terms of "\pi"

2) When we are mentioning "volume is decreasing" then we don't need to give negative sign and so it (answer) will be a positive number. But when we are writing in mathematical form "\frac{dV}{dt}" , we need to give -ve sign to show that it is decreasing.